# Solar Cruiser Disturbance Torque Estimation and Predictive Momentum Management

### v1

@@ 2.3.1 Attitude Control Law
${\boldsymbol{\tau}}^{\text{RW}}_{b,\text{des}}=-\dot{\mbf{h}}^{\text{RW}}_{b,\text{des}}=-\mbf{K}_{p}{\tilde{\boldsymbol{\theta}}}(t)-\mbf{K}_{d}\dot{{\tilde{\boldsymbol{\theta}}}}(t)-\mbf{K}_{i}\int^{t}_{t_{0}}{\tilde{\boldsymbol{\theta}}}(\tau)\textrm{d}\tau,$ (2)
where ${\tilde{\boldsymbol{\theta}}}(t)={\boldsymbol{\theta}}(t)-{\boldsymbol{\theta}}_{d}$, $\dot{{\tilde{\boldsymbol{\theta}}}}(t)=\dot{{\boldsymbol{\theta}}}(t)-\dot{{\boldsymbol{\theta}}}_{d}$, and ${\boldsymbol{\theta}}_{d}$, $\dot{{\boldsymbol{\theta}}}_{d}$ are the desired Euler angles and Euler angle rates of the desired trajectory, respectively. In this work, ${\boldsymbol{\theta}}_{d}$ and $\dot{{\boldsymbol{\theta}}}_{d}$ are chosen to be $\mbf{0}$ for simplicity, although this is not a fundamental limitation of the proposed momentum management approach.
@@ 2.4 Momentum Management Control Actuation
The momentum management time step is chosen based on the AMT position command update period of $\Delta t=100$ seconds used on NEA Scout (Orphee et al., 2018), which is much longer than the attitude control (or ADCS) time step that is assumed to be one second in this work. Between momentum management time steps, the AMT and RCD inputs are modeled with zeroth-order-hold (ZOH) discretization. This ZOH is also used when implementing the momentum management and attitude control inputs within the continuous-time numerical simulation of the sailcraft’s nonlinear dynamics.
The AMT actuation input $\mbf{u}^{\text{AMT}}$ is associated with the first two components of $\mbf{r}^{ps}_{b}$, where $\mbf{u}^{\text{AMT}}=\begin{bmatrix}r^{\text{AMT}}_{b1}&r^{\text{AMT}}_{b2}\end{bmatrix}^{\mathsf{T}}$. The layout of Solar Cruiser’s RCDs allows for an approximately pure on-off roll momentum management torque to be generated in either direction (Inness et al., 2023; Tyler et al., 2023; Heaton et al., 2023). Thus, the RCD torque is modeled as ${\boldsymbol{\tau}}_{b}^{\text{RCD}}=\begin{bmatrix}0&0&\tau^{\text{RCD}}_{b3}\end{bmatrix}^{\mathsf{T}}$, where $u^{\text{RCD}}=\tau^{\text{RCD}}_{b3}$ is chosen based on the momentum management strategy. Considering the RCD on-off actuation as an explicit constraint in the MPC optimization problem leads to a mixed-integer problem, which is computationally expensive, and limits the practicality of onboard real-time MPC. To enable the use of convex optimization solvers with the proposed MPC policy, the integer constraint is relaxed, and a PWM quantization is applied to the RCD actuation (Shen and Caverly, 2025b). A continuous value of $u^{\text{RCD}}_{\text{mpc}}$ is allowed in the optimization problem, where $\tau^{\text{RCD}}_{b3,\text{on}}\leq u^{\text{RCD}}_{\text{mpc}}\leq\tau^{\text{RCD}}_{b3,\text{on}}$ and $\tau^{\text{RCD}}_{b3,\text{on}}$ is the roll torque magnitude generated when the RCDs are turned on. After solving the MPC optimization problem, the continuous $u^{\text{RCD}}_{\text{mpc}}$ is then quantized into a discrete value
where $\beta_{\text{on}}\in\{-1,1\}$ denotes the clockwise and counterclockwise directions about the roll ($\underrightarrow{b}^{3}$) axis, $\tau^{\text{RCD}}_{b3,\text{on}}$ denotes the torque magnitude when RCDs are turned on, and $t_{c}=\Delta t\cdot\frac{u^{\text{RCD}}_{\text{mpc}}}{\tau^{\text{RCD}}_{b3,\text{on}}}$ is the length (cut-off time) of a single pulse PWM conversion from a continuous MPC optimal RCD input. In general, $\tau^{\text{RCD}}_{b3,\text{on}}$ depends on the sun incidence angle (SIA), however, a constant SIA, and thus a constant value of $\tau^{\text{RCD}}_{b3,\text{on}}$, is considered in this study. Details of the RCD quantization can be found in the work of Shen and Caverly (2025b).
2.5 Linear Dynamic Model for Estimation and Prediction
@@
The nonlinear dynamic model is linearized about the current state and AMT position at every time step. Since future AMT positions and RCD torques are not known in advance, their rates are assumed to be zero in the linearization. This yields a continuous-time linear time-varying (LTV) model, where the states include the attitude, angular velocity, reaction wheel angular momentum, and an integral state from the integral term of the attitude controller. The state of the linear system is denoted as $\mbf{x}=\begin{bmatrix}{\boldsymbol{\theta}}^{\mathsf{T}}&{\boldsymbol{\omega}}^{ba^{\mathsf{T}}}_{b}&\mbf{h}^{\text{RW}^{\mathsf{T}}}_{b}&\mbf{e}^{\text{int}^{\mathsf{T}}}\end{bmatrix}^{\mathsf{T}}$, where $\mbf{e}^{\text{int}}=\int^{t}_{t_{0}}({\boldsymbol{\theta}}(\tau)-{\boldsymbol{\theta}}_{d})\textrm{d}\tau$ is the internal state representing the integral term of PID law. It is assumed that the SRP force is constant, and perfectly known from the onboard ADCS. External disturbances are represented by $\mbf{w}={\boldsymbol{\tau}}^{\text{dist}}_{b}$, while $\mbf{u}^{\text{AMT}}$ and ${u}^{\text{RCD}}$ denote the AMT position and RCD torque input, respectively. The linearized continuous-time dynamics retaining the first-order term of Taylor series expansion is derived as (Shen and Caverly, 2025b)
$\dot{\mbf{x}}=\mbf{A}\mbf{x}+\mbf{B}_{w}\mbf{w}+\mbf{B}_{u1}\mbf{u}^{\text{AMT}}+\mbf{B}_{u2}u^{\text{RCD}},$ (9)
@@
$\mbf{A}=\mbf{A}(\bar{\mbf{x}},{\bar{\mbf{r}}}^{ps}_{b})=\begin{bmatrix}\mbf{0}_{3\times 3}&\mbf{1}_{3\times 3}&\mbf{0}_{3\times 3}&\mbf{0}_{3\times 3}\\ -{\bar{\mbf{J}}}_{b}^{\mathcal{B}c^{-1}}\mbf{K}_{p}&\frac{\partial\mbf{f}_{2}}{\partial{\boldsymbol{\omega}}_{b}^{ba}}\Bigg|_{\bar{\mbf{x}},\bar{\mbf{u}}}&-{\bar{\mbf{J}}}_{b}^{\mathcal{B}c^{-1}}\bar{{\boldsymbol{\omega}}}_{b}^{ba^{\times}}&-{\bar{\mbf{J}}}_{b}^{\mathcal{B}c^{-1}}\mbf{K}_{i}\\ \mbf{K}_{p}&\mbf{K}_{d}&\mbf{0}_{3\times 3}&\mbf{K}_{i}\\ \mbf{1}_{3\times 3}&\mbf{0}_{3\times 3}&\mbf{0}_{3\times 3}&\mbf{0}_{3\times 3}\end{bmatrix}$,
$\mbf{B}_{w}=\mbf{B}_{w}({\bar{\mbf{r}}}^{ps}_{b})=\begin{bmatrix}\mbf{0}_{3\times 3}\\ {\bar{\mbf{J}}}_{b}^{\mathcal{B}c^{-1}}\\ \mbf{0}_{3\times 3}\\ \mbf{0}_{3\times 3}\end{bmatrix},\hskip 11.49994pt\mbf{B}_{u1}=\mbf{B}_{u1}(\bar{\mbf{x}},{\bar{\mbf{r}}}^{ps}_{b})=\begin{bmatrix}\mbf{0}_{3\times 2}\\ \frac{\partial\mbf{f}_{2}}{\partial\mbf{r}_{b}^{ps}}\Bigg|_{\bar{\mbf{x}},\bar{\mbf{u}}}\begin{bmatrix}\mbf{1}_{2\times 2}\\ \mbf{0}_{1\times 2}\end{bmatrix}\\ \mbf{0}_{3\times 2}\\ \mbf{0}_{3\times 2}\end{bmatrix},\hskip 11.49994pt\mbf{B}_{u2}=\mbf{B}_{u2}({\bar{\mbf{r}}}^{ps}_{b})=\begin{bmatrix}\mbf{0}_{3\times 3}\\ {\bar{\mbf{J}}}_{b}^{\mathcal{B}c^{-1}}\\ \mbf{0}_{3\times 3}\\ \mbf{0}_{3\times 3}\end{bmatrix}\begin{bmatrix}0\\ 0\\ 1\end{bmatrix}$,
@@
$\frac{\partial\mbf{f}_{2}}{\partial{\boldsymbol{\omega}}_{b}^{ba}}\Big|_{\bar{\mbf{x}},{\bar{\mbf{r}}}^{ps}_{b}}$ $={\bar{\mbf{J}}}_{b}^{\mathcal{B}c^{-1}}\Big(\big({\bar{\mbf{J}}}_{b}^{\mathcal{B}c}\bar{{\boldsymbol{\omega}}}_{b}^{ba}\big)^{\times}-\bar{{\boldsymbol{\omega}}}_{b}^{ba^{\times}}{\bar{\mbf{J}}}_{b}^{\mathcal{B}c}+{\bar{\mbf{h}}}^{\text{RWs}^{\times}}_{b}-\mbf{K}_{d}\Big),$
@@
$\frac{\partial\mbf{f}_{2}}{\partial\mbf{r}_{b}^{ps}}\Big|_{\bar{\mbf{x}},{\bar{\mbf{r}}}^{ps}_{b}}$ $={\bar{\mbf{J}}}_{b}^{\mathcal{B}c^{-1}}\Big(-\frac{m_{p}^{3}+m_{s}^{3}}{(m_{p}+m_{s})^{2}}{\bar{\boldsymbol{\omega}}}_{b}^{ba^{\times}}\Big({\bar{\mbf{r}}}^{ps^{\times}}_{b}{\bar{\boldsymbol{\omega}}}_{b}^{ba^{\times}}+\big({\bar{\mbf{r}}}^{ps^{\times}}_{b}{\bar{\boldsymbol{\omega}}}_{b}^{ba}\big)^{\times}\Big)-\frac{m_{p}^{3}+m_{s}^{3}}{(m_{p}+m_{s})^{2}}\Big({\bar{\mbf{r}}}^{ps^{\times}}_{b}\dot{{\bar{\boldsymbol{\omega}}}}_{b}^{ba^{\times}}+\big({\bar{\mbf{r}}}^{ps^{\times}}_{b}\dot{{\bar{\boldsymbol{\omega}}}}_{b}^{ba}\big)^{\times}\Big)-\frac{m_{s}}{m_{p}+m_{s}}\mbf{f}_{b}^{\text{SRP}^{\times}}\Big).$
Note that the state-space matrices depend on ${\bar{\mbf{r}}}^{ps}_{b}$ because $\mbf{J}_{b}^{\mathcal{B}c}$ is a function of $\mbf{r}^{ps}_{b}$. The variables $\bar{\mbf{x}}=\Big[\bar{{\boldsymbol{\theta}}}^{\mathsf{T}}\,\,\,{\bar{{\boldsymbol{\omega}}}^{ba^{\mathsf{T}}}_{b}}\,\,\,\bar{\mbf{h}}^{\text{RWs}^{\mathsf{T}}}_{b}\,\,\,{\bar{\mbf{e}}}^{\text{int}^{\mathsf{T}}}\Big]^{\mathsf{T}}$ and ${\bar{\mbf{r}}}^{ps}_{b}$ represents the current system state and AMT position, respectively, which are chosen as the current values when performing the linearization.
@@
In this work, a ZOH discretization on both AMT and RCD actuation is used, which has a lower computation requirement when compared to the mixed-FOH-ZOH discretization employed by Shen and Caverly (2025a). Compared to a FOH-discretized model, the ZOH-discretization model does not require numerical integration at every momentum management time step. Discretizing Eq. (9) using a ZOH with the momentum management timestep $\Delta t$ results in
$\mbf{x}_{k}=\mbf{A}_{k}\mbf{x}_{k}+\mbf{B}_{w,k}\mbf{w}_{k}+\mbf{B}_{u1,k}\mbf{u}^{\text{AMT}}_{k}+\mbf{B}_{u2,k}{u}^{\text{RCD}}_{k},$ (10)
@@
which is used as the the Kalman filter process model and MPC prediction model. Note that the nonlinear dynamics in Eq. (1) are used as the system’s dynamics for all numerical simulations, while the KF and MPC use the discrete-time LTV model. Section 3 presents a state and disturbance estimation framework based on a Kalman filter that is used to yield the estimates $\hat{\mbf{x}}(t_{k})$ and $\hat{\mbf{w}}(t_{k})$ needed to compute the LTV dynamic model used in MPC.

### v2

@@ 2.3.1 Attitude Control Law
${\boldsymbol{\tau}}^{\text{RW}}_{b,\text{des}}=-\dot{\mbf{h}}^{\text{RW}}_{b,\text{des}}=-\mbf{K}_{p}\bigg({\boldsymbol{\theta}}(t)-{\boldsymbol{\theta}}_{d}(t)\bigg)-\mbf{K}_{d}\bigg({\boldsymbol{\omega}}^{ba}_{b}(t)-{\boldsymbol{\omega}}_{d}(t)\bigg)-\mbf{K}_{i}\int^{t}_{t_{0}}\bigg({\boldsymbol{\theta}}(\tau)-{\boldsymbol{\theta}}_{d}(\tau)\bigg)\textrm{d}\tau,$ (2)
where ${\boldsymbol{\theta}}_{d}(t)$ and ${\boldsymbol{\omega}}_{d}(t)=\mbf{S}^{-1}({\boldsymbol{\theta}}_{d}(t))\dot{{\boldsymbol{\theta}}}_{d}(t)$ are the desired Euler angles and angular velocity of the desired trajectory, respectively. In this work, ${\boldsymbol{\theta}}_{d}$ and $\dot{{\boldsymbol{\theta}}}_{d}$ are chosen to be a smooth trajectory matching a rest-to-rest maneuver with a trapezoidal Euler angle rate profile. Details of the time-dependent desired trajectory can be found in the Appendix. A different trajectory can be selected based on specific mission scenarios.
@@ 2.4 Momentum Management Control Actuation
The momentum management time step is chosen based on the AMT position command update period of $\Delta t=100$ seconds used on NEA Scout (Orphee et al., 2018), which is much longer than the attitude control (or ADCS) time step that is assumed to be one second in this work. Specifically, the AMT and RCD control inputs generated by the momentum management controller are updated at every 100-second momentum management time step, while the numerical simulation of the sailcraft’s nonlinear dynamics runs at a one-second resolution.
The AMT actuation input $\mbf{u}^{\text{AMT}}$ corresponds to the first two components of $\mbf{r}^{ps}_{b}$, where $\mbf{u}^{\text{AMT}}=\begin{bmatrix}r^{\text{AMT}}_{b1}&r^{\text{AMT}}_{b2}\end{bmatrix}^{\mathsf{T}}$. At any given time $t_{k}$ for momentum management update, the AMT actuation dynamics are modeled based on the current position $\mbf{u}^{\text{AMT}}(t_{k})$ and the incoming momentum management command $\mbf{u}^{\text{AMT}}_{k}$. For $t_{k}\leq t\leq t_{k}+\Delta t$, the system drives the AMT actuator along each axis ($i=1,2$) at its maximum speed $|\dot{u}^{\text{AMT}}_{\text{max},i}|$ toward the target command $u^{\text{AMT}}_{k,i}$. Once the commanded position is reached, the actuator stops and holds its position. These resulting transient position $\mbf{r}^{ps}_{b}(t)$ and velocity $\dot{\mbf{r}}^{ps}_{b}(t)$ are then used in the numerical simulation for the nonlinear system dynamics in Eq. (1).
[…]
$\tau^{\text{RCD}}_{b3,\text{on}}=\frac{6.525\times 10^{-5}}{\cos^{2}(17^{\circ})}\cos^{2}(\text{SIA}).$ (8)
2.5 Linear Dynamics for Estimation and Prediction
@@
The nonlinear dynamics model is linearized about the current state and AMT position at every time step, with knowledge of the attitude-dependent SRP force and estimated disturbance torque. This yields a continuous-time linear time-varying (LTV) model, where the states include the attitude, angular velocity, reaction wheel angular momentum, and an integral state from the integral term of the attitude controller. The state of the linear system is denoted as $\mbf{x}=\begin{bmatrix}{\boldsymbol{\theta}}^{\mathsf{T}}&{\boldsymbol{\omega}}^{ba^{\mathsf{T}}}_{b}&\mbf{h}^{\text{RW}^{\mathsf{T}}}_{b}&\mbf{e}^{\text{int}^{\mathsf{T}}}\end{bmatrix}^{\mathsf{T}}$, where $\mbf{e}^{\text{int}}=\int^{t}_{t_{0}}\big({\boldsymbol{\theta}}(\tau)-{\boldsymbol{\theta}}_{d}(\tau)\big)\textrm{d}\tau$ is the internal state representing the integral term of PID law. It is assumed that the SRP force $\bar{\mbf{f}}_{b}^{\text{SRP}}$ is known through an estimate from the onboard ADCS. External disturbances are represented by $\mbf{w}={\boldsymbol{\tau}}^{\text{dist}}_{b}$, while $\mbf{u}^{\text{AMT}}$ and ${u}^{\text{RCD}}$ denote the AMT position and RCD torque input, respectively. In order to perform time-varying trajectory tracking of ${\boldsymbol{\theta}}_{d}$ and ${\boldsymbol{\omega}}_{d}$, the system is linearized about the operation point of the desired trajectory with the current angular momentum and integral state $\bar{\mbf{x}}=\Big[{\boldsymbol{\theta}}_{d}^{\mathsf{T}}\,\,\,{\boldsymbol{\omega}}_{d}^{{\mathsf{T}}}\,\,\,\bar{\mbf{h}}^{\text{RW}^{\mathsf{T}}}_{b}\,\,\,{\bar{\mbf{e}}}^{\text{int}^{\mathsf{T}}}\Big]^{\mathsf{T}}$, the current AMT position ${\bar{\mbf{r}}}^{ps}_{b}$, a nominally-off RCD torque $\bar{{\boldsymbol{\tau}}}^{\text{RCD}}_{b}=\mbf{0}$, and the current SRP force $\bar{\mbf{f}}_{b}^{\text{SRP}}$, which are chosen as the current values when performing the linearization. Since future AMT positions are not known in advance, their rates are assumed to be zero in the linearization, i.e., $\dot{\bar{\mbf{r}}}_{b}^{ps}=\mbf{0}$, $\ddot{\bar{\mbf{r}}}_{b}^{ps}=\mbf{0}$. The linearized continuous-time dynamics retaining the first-order term of Taylor series expansion is derived as
$\dot{\mbf{x}}=\mbf{A}\mbf{x}+\mbf{B}_{w}\mbf{w}+\mbf{B}_{u1}\mbf{u}^{\text{AMT}}+\mbf{B}_{u2}u^{\text{RCD}}+\mbf{z},$ (9)
@@
$\mbf{A}=\mbf{A}(\bar{\mbf{x}},{\bar{\mbf{r}}}^{ps}_{b})=\begin{bmatrix}\mbf{0}_{3\times 3}&\mbf{1}_{3\times 3}&\mbf{0}_{3\times 3}&\mbf{0}_{3\times 3}\\ -{\bar{\mbf{J}}}_{b}^{\mathcal{B}c^{-1}}\mbf{K}_{p}&\frac{\partial\mbf{f}_{2}}{\partial{\boldsymbol{\omega}}_{b}^{ba}}\Bigg|_{\bar{\mbf{x}},{\bar{\mbf{r}}}^{ps}_{b}}&-{\bar{\mbf{J}}}_{b}^{\mathcal{B}c^{-1}}\bar{{\boldsymbol{\omega}}}_{b}^{ba^{\times}}&-{\bar{\mbf{J}}}_{b}^{\mathcal{B}c^{-1}}\mbf{K}_{i}\\ \mbf{K}_{p}&\mbf{K}_{d}&\mbf{0}_{3\times 3}&\mbf{K}_{i}\\ \mbf{1}_{3\times 3}&\mbf{0}_{3\times 3}&\mbf{0}_{3\times 3}&\mbf{0}_{3\times 3}\end{bmatrix}$,
$\mbf{B}_{w}=\mbf{B}_{w}({\bar{\mbf{r}}}^{ps}_{b})=\begin{bmatrix}\mbf{0}_{3\times 3}\\ {\bar{\mbf{J}}}_{b}^{\mathcal{B}c^{-1}}\\ \mbf{0}_{3\times 3}\\ \mbf{0}_{3\times 3}\end{bmatrix},\hskip 11.49994pt\mbf{B}_{u1}=\mbf{B}_{u1}(\bar{\mbf{x}},{\bar{\mbf{r}}}^{ps}_{b},{\bar{\mbf{f}}}_{b}^{\text{SRP}})=\begin{bmatrix}\mbf{0}_{3\times 2}\\ \frac{\partial\mbf{f}_{2}}{\partial\mbf{r}_{b}^{ps}}\Bigg|_{\bar{\mbf{x}},{\bar{\mbf{r}}}^{ps}_{b},{\bar{\mbf{f}}}_{b}^{\text{SRP}}}\begin{bmatrix}\mbf{1}_{2\times 2}\\ \mbf{0}_{1\times 2}\end{bmatrix}\\ \mbf{0}_{3\times 2}\\ \mbf{0}_{3\times 2}\end{bmatrix},\hskip 11.49994pt\mbf{B}_{u2}=\mbf{B}_{u2}({\bar{\mbf{r}}}^{ps}_{b})=\begin{bmatrix}\mbf{0}_{3\times 3}\\ {\bar{\mbf{J}}}_{b}^{\mathcal{B}c^{-1}}\\ \mbf{0}_{3\times 3}\\ \mbf{0}_{3\times 3}\end{bmatrix}\begin{bmatrix}0\\ 0\\ 1\end{bmatrix}$,
$\mbf{z}=\mbf{z}(\bar{\mbf{x}},{\bar{\mbf{r}}}^{ps}_{b},\bar{{\boldsymbol{\tau}}}^{\text{dist}}_{b},{\bar{\mbf{f}}}_{b}^{\text{SRP}},\bar{{\boldsymbol{\tau}}}_{b}^{\text{dist}})=\mbf{f}(\bar{\mbf{x}},{\bar{\mbf{r}}}^{ps}_{b},\bar{{\boldsymbol{\tau}}}^{\text{RCD}}_{b},{\bar{\mbf{f}}}_{b}^{\text{SRP}},\bar{{\boldsymbol{\tau}}}_{b}^{\text{dist}})-\mbf{A}\bar{\mbf{x}}-\mbf{B}_{w}\bar{{\boldsymbol{\tau}}}^{\text{dist}}_{b}-\mbf{B}_{u1}{\bar{\mbf{r}}}^{ps}_{b},$
@@
$\frac{\partial\mbf{f}_{2}}{\partial{\boldsymbol{\omega}}_{b}^{ba}}\Big|_{\bar{\mbf{x}},{\bar{\mbf{r}}}^{ps}_{b}}={\bar{\mbf{J}}}_{b}^{\mathcal{B}c^{-1}}\Big(\big({\bar{\mbf{J}}}_{b}^{\mathcal{B}c}\bar{{\boldsymbol{\omega}}}_{b}^{ba}\big)^{\times}-\bar{{\boldsymbol{\omega}}}_{b}^{ba^{\times}}{\bar{\mbf{J}}}_{b}^{\mathcal{B}c}+{\bar{\mbf{h}}}^{\text{RW}^{\times}}_{b}-\mbf{K}_{d}\Big),$
@@
$\frac{\partial\mbf{f}_{2}}{\partial\mbf{r}_{b}^{ps}}\Big|_{\bar{\mbf{x}},{\bar{\mbf{r}}}^{ps}_{b},{\bar{\mbf{f}}}_{b}^{\text{SRP}},\bar{{\boldsymbol{\tau}}}_{b}^{\text{dist}}}={\bar{\mbf{J}}}_{b}^{\mathcal{B}c^{-1}}\Big(-\frac{m_{p}^{3}+m_{s}^{3}}{(m_{p}+m_{s})^{2}}{\bar{\boldsymbol{\omega}}}_{b}^{ba^{\times}}\Big({\bar{\mbf{r}}}^{ps^{\times}}_{b}{\bar{\boldsymbol{\omega}}}_{b}^{ba^{\times}}+\big({\bar{\mbf{r}}}^{ps^{\times}}_{b}{\bar{\boldsymbol{\omega}}}_{b}^{ba}\big)^{\times}\Big)-\frac{m_{p}^{3}+m_{s}^{3}}{(m_{p}+m_{s})^{2}}\Big({\bar{\mbf{r}}}^{ps^{\times}}_{b}\dot{{\bar{\boldsymbol{\omega}}}}_{b}^{ba^{\times}}+\big({\bar{\mbf{r}}}^{ps^{\times}}_{b}\dot{{\bar{\boldsymbol{\omega}}}}_{b}^{ba}\big)^{\times}\Big)-\frac{m_{s}}{m_{p}+m_{s}}{\bar{\mbf{f}}}_{b}^{\text{SRP}^{\times}}\Big).$
Note that the state-space matrices depend on ${\bar{\mbf{r}}}^{ps}_{b}$ because $\mbf{J}_{b}^{\mathcal{B}c}$ is a function of $\mbf{r}^{ps}_{b}$. The nonlinear function is defined as
$\mbf{f}(\bar{\mbf{x}},{\bar{\mbf{r}}}^{ps}_{b},\bar{{\boldsymbol{\tau}}}^{\text{RCD}}_{b},{\bar{\mbf{f}}}_{b}^{\text{SRP}},\bar{{\boldsymbol{\tau}}}_{b}^{\text{dist}})=\begin{bmatrix}\mbf{S}^{-1}({\boldsymbol{\theta}}_{d})\\ \mbf{f}_{2}(\bar{\mbf{x}},{\bar{\mbf{r}}}^{ps}_{b},\bar{{\boldsymbol{\tau}}}^{\text{RCD}}_{b},{\bar{\mbf{f}}}_{b}^{\text{SRP}},\bar{{\boldsymbol{\tau}}}_{b}^{\text{dist}})\\ \mbf{K}_{i}\bar{\mbf{e}}^{\text{int}}\\ \mbf{0}\end{bmatrix}$,
where
@@
$\mbf{f}_{2}=\dot{{\boldsymbol{\omega}}}_{b}^{ba}=\mbf{J}_{b}^{{\mathcal{B}c}^{-1}}\Bigg($ $-{\boldsymbol{\omega}}_{b}^{ba^{\times}}\mbf{J}_{b}^{\mathcal{B}c}{\boldsymbol{\omega}}_{b}^{ba}-{\boldsymbol{\omega}}_{b}^{ba^{\times}}\mbf{h}_{b}^{\text{RW}}-\frac{m_{p}^{3}+m_{s}^{3}}{(m_{p}+m_{s})^{2}}\bigg(\mbf{r}_{b}^{ps^{\times}}\ddot{\mbf{r}}_{b}^{ps}-2\dot{\mbf{r}}_{b}^{ps^{\times}}\mbf{r}_{b}^{ps^{\times}}{\boldsymbol{\omega}}_{b}^{ba}\bigg)$
$+\frac{m_{s}}{m_{p}+m_{s}}\mbf{r}_{b}^{ps^{\times}}\mbf{f}_{b}^{\text{SRP}}+{\boldsymbol{\tau}}_{b}^{\text{RCD}}+{\boldsymbol{\tau}}_{b}^{\text{dist}}-\mbf{K}_{p}({\boldsymbol{\theta}}-{\boldsymbol{\theta}}_{d})-\mbf{K}_{d}({\boldsymbol{\omega}}^{ba}_{b}-{\boldsymbol{\omega}}_{d})-\mbf{K}_{i}\int^{t}_{t_{0}}\bigg({\boldsymbol{\theta}}(\tau)-{\boldsymbol{\theta}}_{d}(\tau)\bigg)\textrm{d}\tau\Bigg)$
is a combination of the attitude dynamics in Eq. (1) and the RW control law in Eq. (2) that characterize $\dot{{\boldsymbol{\omega}}}_{b}^{ba}$.
@@
In this work, a zeroth-order hold (ZOH) discretization on both AMT and RCD actuation is used, which has a lower computation requirement when compared to the mixed-FOH-ZOH discretization employed by Shen and Caverly (2025). Discretizing Eq. (9) using a ZOH with the momentum management timestep $\Delta t$ results in
$\mbf{x}_{k}=\mbf{A}_{k}\mbf{x}_{k}+\mbf{B}_{w,k}\mbf{w}_{k}+\mbf{B}_{u1,k}\mbf{u}^{\text{AMT}}_{k}+\mbf{B}_{u2,k}{u}^{\text{RCD}}_{k}+\mbf{z}_{k},$ (10)
which is used as the the Kalman filter process model and MPC prediction model. The discrete-time LTV state-space matrices are obtained by solving $\mbf{A}_{k}={\boldsymbol{\Phi}}(t_{k+1},t_{k})$, $\begin{bmatrix}\mbf{B}_{w,k}&\mbf{B}_{u1,k}&\mbf{B}_{u2,k}\end{bmatrix}=\mbf{A}_{k}\tilde{\mbf{B}}_{k}$, and $\mbf{z}_{k}=\mbf{A}_{k}\tilde{\mbf{z}}_{k}$ through numerical integration of the matrix differential equations
$\dot{{\boldsymbol{\Phi}}}(t,t_{k})=\mbf{A}{\boldsymbol{\Phi}}(t,t_{k}),$
$\dot{\tilde{\mbf{B}}}_{k}(t,t_{k})={\boldsymbol{\Phi}}^{-1}(t,t_{k})\begin{bmatrix}\mbf{B}_{w}&\mbf{B}_{u1}&\mbf{B}_{u2}\end{bmatrix}$,
$\dot{\tilde{\mbf{z}}}_{k}(t,t_{k})={\boldsymbol{\Phi}}^{-1}(t,t_{k})\mbf{z},$
@@
with the initial values ${\boldsymbol{\Phi}}(t_{k+1},t_{k})=\mbf{1}$, $\tilde{\mbf{B}}_{k}(t_{k})=\mbf{0}$, $\tilde{\mbf{z}}_{k}(t_{k})=\mbf{0}$ over the time interval $t\in[t_{k},t_{k+1}]$. Note that the nonlinear dynamics in Eq. (1) are used as the system’s dynamics for all numerical simulations, while the Kalman filter and MPC use the discrete-time LTV model. Section 3 presents a state and disturbance estimation framework based on a Kalman filter that is used to yield the estimates $\hat{\mbf{x}}(t_{k})$ and $\hat{\mbf{w}}(t_{k})$ needed to compute the LTV dynamics model used in MPC.